Identify the element being oxidized and the element being reduced in the following reaction: P4(s)+5O2(g)->P4O10(s)
1
Expert's answer
2012-11-09T05:54:39-0500
P40(s) + 5O20(g)= P4+5O10-2(s)
P40 – 20е = 4P+5 20 4 1 -oxidized
O20 + 4е = 2O-2 4 20 5 - reduced